# New bounds for proper $h$-conflict-free colourings

### v1

@@ 3.2. The proof
By slightly modifying the greedy algorithm presented in the proof of 3, we obtain a colouring which uses $\Delta$ fewer colours, at the cost of granting one fewer witness for vertices of large degree.
@@
Given a graph $G$ of maximum degree $\Delta$, an integer $h\geq 1$, and a proper $k$-colouring $\sigma_{0}$ of $G$ which is $\langle V_{\leq\frac{h}{h+1}\Delta}\;?\;h:h-1\rangle$-conflict-free, as provided by Lemma 18, we want to make small alterations to $\sigma_{0}$ in such a way that
@@
We do this by introducing $m$ additional colours, and thus obtain an $h$-pcf $(k+m)$-colouring. This is possible for some $m=\mathcal{O}\left(h\log\Delta\right)$; this is proved in Lemma 19 using a Rödl Nibble approach.
@@
Let $G$ be a graph of maximum degree $\Delta\geq 30000$. Let $h\geq h_{0}$, $d$ and $\eta$ be non-negative integers such that $h_{0}\leq\frac{d}{4}$ and $28\log\Delta\leq\eta\leq\frac{d}{100}$. Suppose there exists a proper $k$-colouring $\sigma_{0}$ of $G$ that is $\langle V_{\leq d}\;?\;h:h_{0}\rangle$-conflict-free. Let $m\coloneqq\left\lceil 8e(h+2)\Delta\eta/d\right\rceil$. Then there exists a proper $(k+m)$-colouring $\sigma$ of $G$ that is $\langle V_{\leq d}\;?\;h:h_{0}+\eta\rangle$-conflict-free.
@@
We define $p\coloneqq 1/(h+2)\Delta$ and execute the following procedure, which first initialises subsets $A_{1}\dots,A_{m}$ by randomly sampling $V(G)$ with independent probability $p$, and then proceeds in $m$ rounds to construct subsets $C_{1},\dots,C_{m}$ which will be used to modify the colouring $\sigma_{0}$.
@@
       $S_{i}\leftarrow\{u\in V(G):\mathcal{N}(u)\cap A_{i}\neq\emptyset\}\ \cup\ \bigcup_{j=i_{0}}^{i-1}C_{j}$; $C_{i}\leftarrow A_{i}\setminus S_{i}$.
@@
For $i\in[m]$, we respectively let $\mathbf{A}_{i}$, $\mathbf{S}_{i}$, $\mathbf{C}_{i}$, and $\mathbf{i_{0}}$ be the random values of the variables $A_{i}$, $S_{i}$, $C_{i}$, and $i_{0}$ at the end of the procedure. Let ${\bm{\sigma}}$ be the random colouring obtained from $\sigma_{0}$ after recolouring $\mathbf{C}_{i}$ with the colour $k+i$ for each $i\in\{\mathbf{i_{0}},\dots,m\}$. Note that there is no ambiguity in the definition of ${\bm{\sigma}}$ since the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$ are disjoint by construction.
@@
Let $\mathbf{Z}_{v}\coloneqq\sum_{i=1}^{m}|\mathbf{A}_{i}\cap N(v)|\sim B(m\deg(v),p)$. We have $2m\deg(v)p<\frac{\deg(v)}{2}$ by the assumption $\eta\leq\frac{d}{100}$, therefore,
$\mathbb{P}\left[B_{v}\right]\leq\mathbb{P}\left[\mathbf{Z}_{v}\geq\frac{\deg(v)}{2}\right]\leq\mathbb{P}\left[B\left(m\deg(v),p\right)\geq 2m\deg(v)p\right]\;\leq\;e^{-m\deg(v)p/3}<\Delta^{-7},$
where we have used Lemma 13(ii) together with the fact that $m\deg(v)p/3\geq mdp/3\geq 8e\eta/3>7\log\Delta$. ∎
@@
For $v\in V_{>d}$, let $\mathbf{X}_{v}^{i}$ be the indicator variable for the event “$N(v)\cap\mathbf{C}_{i}=\{u\}$ for some $u\notin\mathcal{W}(v)$”. The following fact is easily deduced from the disjointness of the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$ and the previous fact.
@@
Let $\mathbf{X}_{v}\coloneqq\sum_{i=1}^{m}\mathbf{X}_{v}^{i}$. If $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$, then $v$ has at least $h_{0}+\eta$ witnesses in ${\bm{\sigma}}$, as desired. Let $E_{v}$ be the (bad) event that $\mathbf{X}_{v}<\eta$.
@@
Let $v\in V_{>d}$. Then $\mathbb{P}\left[E_{v}\right]\leq\Delta^{-7}$.
@@
Consider $i\in[m]$. Let us fix the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and the resulting values of $\mathbf{C}_{1},\dots,\mathbf{C}_{i-1}$. Let $C^{\prime}$ be the resulting value of the term “$\bigcup_{j=i_{0}}^{i-1}C_{j}$” at the start of round $i$. The distribution of $\mathbf{A}_{i}$ is unaffected by this conditioning, i.e. every vertex $v$ is still present in $\mathbf{A}_{i}$ independently with probability $p$. Let $U\coloneqq N(v)\setminus(\mathcal{W}(v)\cup C^{\prime})$ and let $t\coloneqq|U|$. Because of the fail-safe, $C^{\prime}$ contains at most $\deg(v)/2$ neighbours of $v$, and $|\mathcal{W}(v)|=h_{0}\leq\frac{d}{4}$, therefore $t\geq\deg(v)-h_{0}-\frac{\deg(v)}{2}\geq\frac{d}{4}$.
The probability that $N(v)\cap\mathbf{A}_{i}=\{u\}$ with $u\in U$ is equal to $tp(1-p)^{\deg(v)-1}\geq\frac{dp}{4}(1-p)^{\Delta-1}$. Suppose such a vertex $u\in U$ exists, i.e. we are conditioning on the event “$N(v)\cap\mathbf{A}_{i}=\{u\}$ for some $u\in U$”. By independence of the random choices, the distribution of $\mathbf{A}_{i}$ outside of $N(v)$ remains unaffected, i.e. every vertex $x\in V\setminus N(v)$ is present in $\mathbf{A}_{i}$ independently with probability $p$. We now determine a lower bound on the probability that $u\in\mathbf{C}_{i}$, which is equivalent to $u\notin\mathbf{S}_{i}$ since $u\in\mathbf{A}_{i}$. By definition of $U$, $u\notin C^{\prime}$, therefore $u\notin\mathbf{S}_{i}$ if and only if $\mathcal{N}(u)\cap\mathbf{A}_{i}=\emptyset$: we already know that $N(v)\cap\mathbf{A}_{i}=\{u\}$, so this happens with probability $(1-p)^{|\mathcal{N}(u)\setminus N(v)|}\geq(1-p)^{(h+1)\Delta}$. Therefore,
$\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]\geq\frac{dp}{4}(1-p)^{(h+2)\Delta-1}>\frac{dp}{4e},$ (1)
where we have used the fact that $(1-\frac{1}{x})^{x-1}>\frac{1}{e}$ for all $x\geq 1$. Let $q\coloneqq\frac{dp}{4e}$. We have shown that $\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]>q$ regardless of the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and so in particular,
@@
By Lemma 15, we obtain $\mathbb{P}\left[\sum\mathbf{X}_{v}^{i}<\eta\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]$. We have defined $m$ such that $mq\geq 2\eta$, and since $\eta\geq 28\log\Delta$ we have $\frac{mq}{8}\geq 7\log\Delta$; by Lemma 13(i), we conclude that $\mathbb{P}\left[E_{v}\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]\leq\Delta^{-7}$. ∎
For $v\in V_{>d}$, the bad events $B_{v}$ and $E_{v}$ are determined by the realisation of $(\mathbf{A}_{1},\ldots,\mathbf{A}_{m})$ within the distance-$3$ neighbourhood of $v$, therefore these bad events are independent from all but at most $2\Delta^{6}$ other such bad events. By Lemma 16, with nonzero probability, none of the events $B_{v}$ or $E_{v}$ occur. In that case, $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$ for every $v\in V_{>d}$, so by 21 and 22, we conclude that the colouring ${\bm{\sigma}}$ is proper and $\langle V_{\leq d}\;?\;h:h_{0}+\eta\rangle$-conflict-free. ∎
5 is a direct consequence of the following corollary, which provides explicit constants.
@@
We may assume that $h\leq\frac{\Delta}{8}$, since otherwise $700(h+5)\log\Delta>\Delta$ and we already know that $\chi_{\rm pcf}^{h}(G)\leq(h+1)\Delta+1$ by 3.
By Lemma 18, there exists a proper $\langle V_{\leq\frac{h}{h+1}\Delta}\;?\;h:h-1\rangle$-conflict-free $(h\Delta+1)$-colouring $\sigma_{0}$ of $G$. We apply Lemma 19 with $k\coloneqq h\Delta+1$, $d\coloneqq\frac{h}{h+1}\Delta$, $h_{0}\coloneqq h-1$, and $\eta\coloneqq\left\lceil 28\log\Delta\right\rceil$. We have $h_{0}\leq d/4$ by the assumption $h\leq\frac{\Delta}{8}$, and we have $\eta\leq\frac{d}{100}$ by the assumption $\Delta\geq 70000$. Then, we obtain a proper $(k+m)$-colouring $\sigma$ that is $\langle V_{\leq d}\;?\;h:h-1+\eta\rangle$-conflict-free, where $m=\left\lceil 8e(h+2)\Delta\eta/d\right\rceil$. A quick computation shows that $k+m\leq h\Delta+700(h+5)\log\Delta$. ∎
@@ 4. Bounding $\chi_{\rm pcf}^{h}$ for graphs of constrained minimum degree
The first-order term $h\Delta$ of the upper bound in 5 is tight, by 6: the lower bound $h\Delta+1$ is certified by a graph construction of minimum degree $h+1$. In fact, such graphs are highly constrained due to the requirement that neighbourhoods of size at most $h+1$ must be rainbow in an $h$-pcf colouring. If the minimum degree $\delta$ is sufficiently large with respect to $h$ and $\Delta$, then one would expect more flexibility when dealing with small neighbourhoods. Indeed the upper bound on $\chi_{\rm pcf}^{h}(G)$ from 9 established by Kamyczura and Przybyło in [9] shows that only a little more than $\Delta$ colours are required for an $h$-pcf colouring in that case, with the extra assumption that $\delta$ is sufficiently large with respect to $\Delta$.
We wish to drop the latter extra assumption, and consider a minimum degree constraint as weak as possible that still ensures that $\chi_{\rm pcf}^{h}$ is $(1+o(1))\Delta$. To do so, we again rely on a two-step colouring process, that takes advantage of having large minimum degree. In the previous section, the first step described by Lemma 18 provides exactly $h$ witnesses for vertices of small degree; as a consequence, during the second step described in Lemma 19, we need to ensure that no witness created during the first step is destroyed, which is very costly. In contrast, the extra minimum degree condition lets us replace the deterministic greedy first step with a random one that creates more witnesses for vertices of small degree, namely $2h$ instead of $h$, so that during the second step we may destroy a fraction of those witnesses. For that, we require that $\delta\geq 16h$, and in Lemma 25 we show that a uniformly random proper colouring with sufficiently many colours satisfies the required conditions with non-zero probability.
The strategy for the second step is essentially the same as in Section 3; we still rely on a Rödl Nibble approach to recolour a fraction of the vertices in such a way that new witnesses are created for large-degree vertices. The fact that we are allowed to destroy some witnesses from the precolouring makes this second step much cheaper than the one from the previous section. This is presented in Lemma 27.
@@
[…]
@@
Let $k\coloneqq\Delta+3d$, and let $\mathscr{C}_{k}(G)$ be the set of all proper $k$-colourings of $G$. Given a proper colouring $\sigma\in\mathscr{C}_{k}(G)$ and a vertex $v\in V(G)$, we let $L_{\sigma}(v)\coloneqq[k]\setminus\sigma(N(v))$ be the set of colours $x$ such that, if we redefine $\sigma(v)\leftarrow x$, we still have $\sigma\in\mathscr{C}_{k}(G)$.
[…]
For every $i\leq\deg(v)$, let $\mathbf{S}_{i}$ be the number of witnesses of $v$ in ${\bm{\sigma}}_{i}$. For every $i\geq 1$, we have $\mathbf{S}_{i}=\mathbf{S}_{i-1}+\mathbf{X}_{i}$ where $\mathbf{X}_{i}=-1$ if ${\bm{\sigma}}_{i}(u_{i})$ is the colour of one of the $\mathbf{S}_{i-1}$ witnesses of $v$ in ${\bm{\sigma}}_{i-1}$; $\mathbf{X}_{i}=1$ if ${\bm{\sigma}}_{i}(u_{i})\notin{\bm{\sigma}}_{i-1}(N(v))$; and $\mathbf{X}_{i}=0$ otherwise. Let $\sigma_{i-1}$ be any possible realisation of ${\bm{\sigma}}_{i-1}$. Since there are always at least $3d$ colours available when resampling the colour of $u_{i}$, and at most $d$ colours that appear in $N(v)$, we have
@@
We now condition on ${\bm{\sigma}}=\sigma_{0}$, and prove that under that condition $\mathbf{S}_{\deg(v)}\geq 2h$ with large enough probability. Note that under that condition, the realisation of ${{\bm{\sigma}}_{\deg(v)}}_{\left|V(G)\setminus N(v)\right.}$ is precisely ${\sigma_{0}}_{\left|V(G)\setminus N(v)\right.}$. Let $\{\mathbf{Y}_{i}\}$ be a sequence of Bernoulli random variables such that
@@
Hence by Lemma 15 we have $\mathbb{P}\left[\sum\mathbf{Y}_{i}\leq t\right]\leq\mathbb{P}\left[B(\deg(v),2/3)\leq t\right]$. If $B_{{\bm{\sigma}}}(v)$ occurs, then it must hold that $\sum\mathbf{Y}_{i}<\deg(v)/2+h$, and since $h\leq\delta/16$, then in particular it must also hold that $\sum\mathbf{Y}_{i}<\frac{9}{16}\deg(v)=(1-5/32)\frac{2}{3}\deg(v)$. With an application of Chernoff’s bound, we obtain

### v2

@@ 3.2. The proof
By slightly modifying the greedy algorithm presented in the proof of ˜3, we obtain a colouring which uses $\Delta$ fewer colours, at the cost of granting one fewer witness for vertices of large degree.
@@
Given a graph $G$ of maximum degree $\Delta$, an integer $h\geq 1$, and a proper $k$-colouring $\sigma_{0}$ of $G$ which is $\langle V_{\leq\frac{h}{h+1}\Delta}\;?\;h:h-1\rangle$-conflict-free, as provided by Lemma˜18, we want to make small alterations to $\sigma_{0}$ in such a way that
@@
We do this by introducing $m$ additional colours, and thus obtain an $h$-pcf $(k+m)$-colouring. This is possible for some $m=\mathcal{O}\left(h\log\Delta\right)$; this is proved in Lemma˜19 using a Rödl Nibble approach.
@@
Let $G$ be a graph of maximum degree $\Delta\geq 13000$. Let $h$, $h_{0}$, $d$, $\eta$ be non-negative integers such that $h_{0}\leq\frac{d}{4}$ and $25\log\Delta\leq\eta\leq\frac{d}{20e}$. Suppose there exists a proper $k$-colouring $\sigma_{0}$ of $G$ that is $\langle V_{\leq d}\;?\;h:h_{0}\rangle$-conflict-free. Let $m\coloneqq\left\lceil 4e(h+2)\Delta\eta/d\right\rceil$. Then there exists a proper $(k+m)$-colouring $\sigma$ of $G$ that is $\langle V_{\leq d}\;?\;h:h_{0}+\eta\rangle$-conflict-free.
@@
We define $p\coloneqq\frac{1}{(h+2)\Delta}$ and execute the following procedure, which first initialises subsets $A_{1}\dots,A_{m}$ by randomly sampling $V(G)$ with independent probability $p$, and then proceeds in $m$ rounds to construct subsets $C_{1},\dots,C_{m}$ which will serve as new colour classes in the colouring $\sigma_{0}$. We note that the randomness of the procedure is entirely contained in the Initialisation bloc, which simply constructs $m$ independent binomial random subsets of $V(G)$. The fail-safe bloc could be replaced with an interruption of the procedure, but its presence makes the analysis easier by preserving some crucial invariants needed in the rest of the proof.
@@
 $S_{i}\leftarrow\{u\in V(G):\mathcal{N}(u)\cap A_{i}\neq\emptyset\}\ \cup\ \bigcup_{j=i_{0}}^{i-1}C_{j}$
 $C_{i}\leftarrow A_{i}\setminus S_{i}$.
 fail-safe: if there exists $v\in V_{>d}$ such that $|N(v)\cap\bigcup_{j=i_{0}}^{i}C_{j}|>\frac{\deg(v)}{4}$ then
@@
For $i\in[m]$, we respectively let $\mathbf{A}_{i}$, $\mathbf{S}_{i}$, $\mathbf{C}_{i}$, and $\mathbf{i_{0}}$ be the random values of the sets $A_{i}$, $S_{i}$, $C_{i}$ and the integer $i_{0}$ at the end of the procedure. Let $\bm{\sigma}$} be the random colouring obtained from $\sigma_{0}$ after recolouring $\mathbf{C}_{i}$ with the colour $k+i$ for each $i\in\{\mathbf{i_{0}},\dots,m\}$. Note that there is no ambiguity in the definition of $\bm{\sigma}$} since the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$ are disjoint by construction.
@@
Let $\mathbf{Z}_{v}\coloneqq\sum_{i=1}^{m}|\mathbf{A}_{i}\cap N(v)|\sim B(m\deg(v),p)$. We have $\frac{5}{4}m\deg(v)p<\frac{\deg(v)}{4}$ by the assumption $\eta\leq\frac{d}{20e}$, therefore,
$\mathbb{P}\left[B_{v}\right]\leq\mathbb{P}\left[\mathbf{Z}_{v}\geq\frac{\deg(v)}{4}\right]\leq\mathbb{P}\left[B\left(m\deg(v),p\right)\geq\frac{5}{4}m\deg(v)p\right]\;\leq\;e^{-m\deg(v)p/36}<\Delta^{-7},$
where we have used Lemma˜13(ii) together with the fact that $m\deg(v)p/36\geq mdp/36\geq e\eta/9>7\log\Delta$. ∎
@@
For $v\in V_{>d}$, let $\mathbf{X}_{v}^{i}$ be the indicator variable for the event $N(v)\cap\mathbf{C}_{i}=\{u\}$ for some $u\notin\mathcal{W}(v)$”. The following fact is easily deduced from the disjointness of the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$ and the previous fact.
@@
For every $v\in V_{>d}$, let $\mathbf{X}_{v}\coloneqq\sum_{i=1}^{m}\mathbf{X}_{v}^{i}$. If $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$, then $v$ has at least $h_{0}+\eta$ witnesses in $\bm{\sigma}$}, as desired. Let $E_{v}$ be the (bad) event that $\mathbf{X}_{v}<\eta$.
@@
Let $v\in V_{>d}$. Then $\mathbb{P}\left[E_{v}\right]\leq\Delta^{-6.25}$.
@@
Consider $i\in[m]$. Let us fix the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and the resulting values of $\mathbf{C}_{1},\dots,\mathbf{C}_{i-1}$. Let $C^{\prime}$ be the resulting value of the term $\bigcup_{j=i_{0}}^{i-1}C_{j}$” at the start of round $i$. The distribution of $\mathbf{A}_{i}$ is unaffected by this conditioning, i.e. every vertex $u\in V(G)$ is still present in $\mathbf{A}_{i}$ independently with probability $p$. Let $U\coloneqq N(v)\setminus(\mathcal{W}(v)\cup C^{\prime})$ and let $t\coloneqq|U|$. Because of the fail-safe, $C^{\prime}$ contains at most $\deg(v)/4$ neighbours of $v$, and $|\mathcal{W}(v)|=h_{0}\leq\frac{d}{4}$, therefore $t\geq\deg(v)-h_{0}-\frac{\deg(v)}{4}\geq\frac{d}{2}$.
The probability that $N(v)\cap\mathbf{A}_{i}=\{u\}$ with $u\in U$ is equal to $tp(1-p)^{\deg(v)-1}\geq\frac{dp}{2}(1-p)^{\Delta-1}$. Suppose such a vertex $u\in U$ exists, i.e. we are conditioning on the event $N(v)\cap\mathbf{A}_{i}=\{u\}$ for some $u\in U$”. By independence of the random choices, the distribution of $\mathbf{A}_{i}$ outside of $N(v)$ remains unaffected, i.e. every vertex $x\in V\setminus N(v)$ is present in $\mathbf{A}_{i}$ independently with probability $p$. We now determine a lower bound on the probability that $u\in\mathbf{C}_{i}$, which is equivalent to $u\notin\mathbf{S}_{i}$ since $u\in\mathbf{A}_{i}$. By definition of $U$, $u\notin C^{\prime}$, therefore $u\notin\mathbf{S}_{i}$ if and only if $\mathcal{N}(u)\cap\mathbf{A}_{i}=\emptyset$. We already know that $N(v)\cap\mathbf{A}_{i}=\{u\}$, so this happens with probability $(1-p)^{|\mathcal{N}(u)\setminus N(v)|}\geq(1-p)^{(h+1)\Delta}$. Therefore,
$\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]\geq\frac{dp}{2}(1-p)^{(h+2)\Delta-1}>\frac{dp}{2e},$ (1)
where we have used the fact that $(1-\frac{1}{x})^{x-1}>\frac{1}{e}$ for all $x>1$. Let $q\coloneqq\frac{dp}{2e}$. We have shown that $\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]>q$ regardless of the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and so in particular,
@@
By Lemma˜15, we obtain $\mathbb{P}\left[\sum\mathbf{X}_{v}^{i}<\eta\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]$. We have defined $m$ such that $mq\geq 2\eta$, and since $\eta\geq 25\log\Delta$ we have $\frac{mq}{8}\geq 6.25\log\Delta$; by Lemma˜13(i), we conclude that $\mathbb{P}\left[E_{v}\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]\leq\Delta^{-6.25}$. ∎
For $v\in V_{>d}$, the bad events $B_{v}$ and $E_{v}$ are determined by the realisation of $(\mathbf{A}_{1},\ldots,\mathbf{A}_{m})$ within the distance-$3$ neighbourhood of $v$, therefore these bad events are independent from all but at most $2\Delta^{6}$ other such bad events. By Lemma˜16, with nonzero probability, none of the events $B_{v}$ or $E_{v}$ occur. In that case, $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$ for every $v\in V_{>d}$, so by ˜21 and ˜22, we conclude that the colouring $\bm{\sigma}$} is proper and $\langle V_{\leq d}\;?\;h:h_{0}+\eta\rangle$-conflict-free. ∎
˜5 is a direct consequence of the following corollary, which provides explicit constants.
@@
We may assume that $h\leq\frac{\Delta}{8}$, since otherwise $100e(h+5)\log\Delta>\Delta$ and we already know that $\chi_{\rm pcf}^{h}(G)\leq(h+1)\Delta+1$ by ˜3.
By Lemma˜18, there exists a proper $\langle V_{\leq\frac{h}{h+1}\Delta}\;?\;h:h-1\rangle$-conflict-free $(h\Delta+1)$-colouring $\sigma_{0}$ of $G$. We apply Lemma˜19 with $k\coloneqq h\Delta+1$, $d\coloneqq\frac{h}{h+1}\Delta\geq\frac{\Delta}{2}$, $h_{0}\coloneqq h-1$, and $\eta\coloneqq\left\lceil 25\log\Delta\right\rceil$. We have $h_{0}\leq d/4$ by the assumption $h\leq\frac{\Delta}{8}$, and we have $\eta\leq\frac{\Delta}{40e}\leq\frac{d}{20e}$ by the assumption $\Delta\geq 28000$. Then, we obtain a proper $(k+m)$-colouring $\sigma$ that is $\langle V_{\leq d}\;?\;h:h-1+\eta\rangle$-conflict-free, where $m=\left\lceil 4e(h+2)\Delta\eta/d\right\rceil$. A quick computation shows that $k+m\leq h\Delta+100e(h+5)\log\Delta$. ∎
@@ 4. Bounding $\chi_{\rm pcf}^{h}$ for graphs of constrained minimum degree
The first-order term $h\Delta$ of the upper bound in ˜5 is tight, by ˜6: the lower bound $h\Delta+1$ is certified by a graph construction of minimum degree $h+1$. In fact, such graphs are highly constrained due to the requirement that neighbourhoods of size at most $h+1$ must be rainbow in an $h$-pcf colouring. If the minimum degree $\delta$ is sufficiently large with respect to $h$ and $\Delta$, then one would expect more flexibility when dealing with small neighbourhoods. Indeed the upper bound on $\chi_{\rm pcf}^{h}(G)$ from ˜9 established by Kamyczura and Przybyło in [10] shows that only a little more than $\Delta$ colours are required for an $h$-pcf colouring in that case, with the extra assumption that $\delta$ is sufficiently large with respect to $\Delta$.
We wish to drop the latter extra assumption, and consider a minimum degree constraint as weak as possible that still ensures that $\chi_{\rm pcf}^{h}$ is $(1+o(1))\Delta$. To do so, we again rely on a two-step colouring process, that takes advantage of having large minimum degree. In the previous section, the first step described by Lemma˜18 provides exactly $h$ witnesses for vertices of small degree; as a consequence, during the second step described in Lemma˜19, we need to ensure that no witness created during the first step is destroyed, which is very costly. In contrast, the extra minimum degree condition lets us replace the deterministic greedy first step with a random one that creates more witnesses for vertices of small degree, namely $2h$ instead of $h$, so that during the second step we may destroy a fraction of those witnesses. For that, we require that $\delta\geq 16h$, and in Lemma˜25 we show that a uniformly random proper colouring with sufficiently many colours satisfies the required conditions with non-zero probability.
The strategy for the second step is essentially the same as in Section 3; we still rely on a Rödl Nibble approach to recolour a fraction of the vertices in such a way that new witnesses are created for large-degree vertices. The fact that we are allowed to destroy some witnesses from the precolouring makes this second step much cheaper than the one from the previous section. This is presented in Lemma˜27.
@@
Let $h\geq 20\log\Delta$ be an integer, and $G$ be a graph of maximum degree $\Delta$ and minimum degree $\delta\geq 16h$. Then, for every $d$ with $\delta\leq d\leq\Delta$, $G$ admits a proper $\langle V_{\leq d}\;?\;2h:0\rangle$-conflict-free $(\Delta+3d)$-colouring.
@@
Let $k\coloneqq\Delta+3d$, and let $\mathscr{C}_{k}(G)$ be the set of all proper $k$-colourings of $G$. Given a proper colouring $\sigma\in\mathscr{C}_{k}(G)$ and a vertex $v\in V(G)$, we let $L_{\sigma}(v)\coloneqq[k]\setminus\sigma(N(v))$ be the set of colours $x$ such that, if we redefine $\sigma(v)\leftarrow x$, we still have $\sigma\in\mathscr{C}_{k}(G)$. We also denote $\sigma\setminus X$ the restriction of $\sigma$ to $V(G)\setminus X$, for every $X\subseteq V(G)$.
We fix some vertex $v\in V_{\leq d}$, and we write $N(v)\coloneqq\{u_{1},u_{2},\ldots,u_{\deg(v)}\}$. Let ${\bm{\sigma}}\in\mathscr{C}_{k}(G)$ be drawn uniformly at random; we want to analyse the probability of the bad event $B_{{\bm{\sigma}}}(v)$ that $v$ has fewer than $2h$ solitary colours in $\bm{\sigma}$}. To do that, we define the sequence of random colourings ${\bm{\sigma}}_{0},\ldots,{\bm{\sigma}}_{\deg(v)}\in\mathscr{C}_{k}(G)$ where ${\bm{\sigma}}_{0}\coloneqq{\bm{\sigma}}$, and for each $i\geq 1$, ${\bm{\sigma}}_{i}$ is obtained from ${\bm{\sigma}}_{i-1}$ be resampling ${\bm{\sigma}}_{i-1}(u_{i})$ uniformly at random from $L_{{\bm{\sigma}}_{i-1}}(u_{i})$. This does not change the probability distribution of the random colourings; each ${\bm{\sigma}}_{i}$ is uniformly distributed in $\mathscr{C}_{k}(G)$.
For every integer $0\leq i\leq\deg(v)$, let $\mathbf{S}_{i}$ be the number of colours that appear exactly once on $\{u_{1},\ldots,u_{i}\}$ in ${\bm{\sigma}}_{i}$. Hence $\mathbf{S}_{\deg(v)}$ is the number of witnesses of $v$ in ${\bm{\sigma}}_{\deg(v)}$. For every $i\in[\deg(v)]$, let $\mathbf{X}_{i}\coloneqq\mathbf{S}_{i}-\mathbf{S}_{i-1}$. Clearly we have that $\mathbf{X}_{i}=-1$ if ${\bm{\sigma}}_{i}(u_{i})$ is one of the colours counted by $\mathbf{S}_{i-1}$; $\mathbf{X}_{i}=1$ if ${\bm{\sigma}}_{i}(u_{i})\notin{\bm{\sigma}}_{i-1}(\{u_{1},\ldots,u_{i-1}\})$; and $\mathbf{X}_{i}=0$ otherwise. Let $\sigma_{i-1}$ be any possible realisation of ${\bm{\sigma}}_{i-1}$. Since there are always at least $3d$ colours available when resampling the colour of $u_{i}$, and at most $d$ colours that appear in $N(v)$, we have
@@
We now condition on ${\bm{\sigma}}=\sigma_{0}$, and prove that under that condition $\mathbf{S}_{\deg(v)}\geq 2h$ with large enough probability. Note that under that condition, the realisation of ${\bm{\sigma}}_{\deg(v)}\setminus N(v)$ is precisely $\sigma_{0}\setminus N(v)$. Let $\{\mathbf{Y}_{i}\}$ be a sequence of Bernoulli random variables such that
@@
Hence by Lemma˜15 we have $\mathbb{P}\left[\sum\mathbf{Y}_{i}\leq t\right]\leq\mathbb{P}\left[B(\deg(v),2/3)\leq t\right]$. If $B_{{\bm{\sigma}}}(v)$ occurs, then it must hold that $\sum\mathbf{Y}_{i}<\deg(v)/2+h$, and since $h\leq\delta/16$, then in particular it must also hold that $\sum\mathbf{Y}_{i}<\frac{9}{16}\deg(v)=(1-5/32)\frac{2}{3}\deg(v)$. With an application of Chernoff’s bound, we obtain